The eyepiece of an astronomical telescope has a focal length of 10 cm. The telescope is focused for normal vision of distant objects when the tube length is 1.0 m. Find the focal length of the objective and the magnifying power of the telescope.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. For the given astronomical telescope in normal adjustment, F base e = 10 cm, L = 1 m = 100cm S0, f base 0 = L – f base e = 100 – 10 = 90 cm and, magnifying power = f base 0/f base e = 90/10 = 9